OpenStudy (anonymous):

Meg has a can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of orange juice and water in the can. Define the variable used in the equation. Part B: How many liters of orange juice are present in the can? Show your work.

OpenStudy (ineedhelplz):

Let x be 80% is orange juice. That means 20% is water. Amount of water = x * 20 /100 = 0.2x The amount of water = 1 liter. 0.2x = 1 is the equation in one variable where x is the total number of liters of orange juice and water combined.

OpenStudy (ineedhelplz):

Part A) 0.2x = 1 where x is the total number of liters of orange juice and water combined. Part B) 0.2x = 1 Solve for x. Orange juice is 80%. The amount of orange juice is: x * 80 / 100. Put x from solving the above equation.

OpenStudy (ineedhelplz):

0.2x = 1 Divide both sides by 0.2 x = 1 / 0.2 = ?

OpenStudy (ineedhelplz):

x = 5 which is the total number of liters of orange juice and water combined. 80% of it is orange juice. Therefore, the number of liters of OJ is: 5 * 80 / 100 = ?

OpenStudy (ineedhelplz):

All you have to do is multiply 5 and 80 and divide by 100.

OpenStudy (ineedhelplz):

5 * 80 / 100 = 4 Part B) 4 liters of orange juice.